OpenStudy (anonymous):

find two positive real numbers whose product is a maximum and whose sum is 154. a)75,79 b)77,77 c)82,72 d)86,68 e)65,89

OpenStudy (anonymous):

hii

OpenStudy (dumbcow):

A + B = 154 A = 154 - B A*B = (154-B)*B = -B^2 + 154B Use formula for finding line of symmetry of this parabola: B = -154/-2 = 77 This value for B maximizes A*B. By substitution A = 154-77 =77 Therefore A = 77, B = 77